OpenStudy (anonymous):

what is 3-7x + 2x when x=3 will fan and medal

OpenStudy (anonymous):

@TIMDAVE

OpenStudy (anonymous):

3-7(3)+2(3) 7(3)= 21 2(3)= 6 3-21+6= -12

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

3 -7(3) + 2(3) = y 3 - 21 + 6 = y 12=y

OpenStudy (anonymous):

-12

OpenStudy (anonymous):

ok thanks so much

OpenStudy (anonymous):

-12 is wrong

OpenStudy (anonymous):

but i will fan u @Anonymous1321

OpenStudy (anonymous):

well my best guess is that -12 could be right

OpenStudy (anonymous):

acutually it was right sorry lol

OpenStudy (anonymous):

ok can i have help again i hate this kinda math